OpenStudy (anonymous):

How to start this proving this equation? cos(3x)=4COS^3x-3COSx

OpenStudy (anonymous):

*How to start proving this equation?

OpenStudy (anonymous):

use the addition angle formula for cosine twice

OpenStudy (anonymous):

\[\cos(3x)=\cos(2x+x)\] for the first step

OpenStudy (anonymous):

Do you mean sum and difference identity for cosine?

OpenStudy (anonymous):

I don't know the addition angle formula. Is there another way of proving this identity?

OpenStudy (raden):

use the formula : cos(A+B) = cos(A)cos(B) - sin(A)sin(B) here, let A=2x and B=x

OpenStudy (anonymous):

what @RadEn said

OpenStudy (anonymous):

then when you get \[\cos(2x)\sin(x)-\sin(2x)\sin(x)\] you can use the double angle formula for \(\sin(2x)\) and \(\cos(2x)\)

OpenStudy (anonymous):

How are you getting 2x?

OpenStudy (raden):

3x=2x+x, right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh. Are ya working from the left side of the problem?

OpenStudy (raden):

yes, but we have to simplify for right side with identities trigono

OpenStudy (raden):

we have : cos(3x) = cos(2x+x) cos(3x) = cos(2x)cos(x) - sin(2x)sin(x)

OpenStudy (raden):

now, use the identities : cos(2x) = 2cos^2 (x) and sin(2x) = 2sin(x)cos(x)

OpenStudy (raden):

so, it can be : cos(3x) = cos(2x+x) = cos(2x)cos(x) - sin(2x)sin(x) = (2cos^2 (x) - 1)cos(x) - 2sin(x)cos(x)sin(x) = 2cos^3 (x) - cos(x) - 2(sin^2 (x))cos(x)

OpenStudy (raden):

so far so good ?

OpenStudy (raden):

sorry, there is typo for identity cos(2x), it should be cos(2x) = 2cos^2 (x) - 1

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, then look the last part still there is sin^2 (x), we have to make all terms in cos, right ?

OpenStudy (anonymous):

right

OpenStudy (raden):

use the identity : sin^2 (x) = 1 - cos^2 (x)

OpenStudy (raden):

so, it can be = .... = 2cos^3 (x) - cos(x) - 2(sin^2 (x))cos(x) = 2cos^3 (x) - cos(x) - 2(1 - cos^2 (x))cos(x) = 2cos^3 (x) - cos(x) - 2cos(x) + 2 cos^3 (x)

OpenStudy (raden):

just combine the similar terms : 2cos^3 (x) + 2cos^3 (x) = 4cos^3 (x) - cos(x) - 2cos(x) = -3cos(x)

OpenStudy (anonymous):

oh wow. thankyou

OpenStudy (raden):

you're welcome :)